[Hallc_running] [New Logentry] RC Update

latif at jlab.org latif at jlab.org
Thu Nov 8 09:20:02 EST 2018

Logentry Text:
- E12-09-017 (P_t scan of SIDIS) finished their last kinematic setting during early swing shift yesterday.
- E12-09-002 (Charge Symmetry Violation) took over during the swing shift after E12-09-017 was over .
- By 8:00 PM yesterday, CSV completed Hydrogen Coincidence Elastics for HMS Checks and went to production.
- Whitney did some test to decide final prescale for singles, and finally concluded the following setting for CSV production:
 PS6 = 0 (HMS and SMHS 3/4 coincidence mode) all the time. Adjust PS4 (HMS) to give at least 100 Hz (and no more than 200 Hz) of HMS singles.
- The beam was very unstable starting around 2:00 AM during the owl shift. MCC informed this morning that they fixed it temporarily and expecting stable beam starting from this morning.
- CSV is currently taking data at kinematic group 2-4.

We still need owl shift takers for coming Saturday, Monday and Tuesday.

The daily run plan is posted here: https://hallcweb.jlab.org/wiki/index.php/RC_daily_meetings_and_daily_runplan_fall_2018

We will meet today at 4:30 PM in the counting house (second floor) to discuss the update and run plan.
Bluejeans link: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3622982
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181108/853373f4/attachment.html>

More information about the Hallc_running mailing list